# Real-time Rendering-based Surgical Instrument Tracking via Evolutionary Optimization

### 12 Mar 2026 v1

@@ I INTRODUCTION
Robot-assisted Minimally-Invasive Surgery (RMIS) has been increasingly adopted to improve surgical safety and procedural efficiency in complex, long-duration operations under physically demanding conditions. Accurate and efficient pose estimation of articulated surgical instruments is essential for a wide range of downstream applications, such as vision-based control [9217086], augmented reality guidance [Kalia2021], and robot learning [kim2024surgicalrobottransformersrt].
Uneven geometric structures, partial visibility of the kinematic chain, and noisy joint angle measurements caused by cable stretch introduce additional challenges for accurate articulated instrument pose recovery. Previous works have approached this problem from both kinematic calibration and pose calibration perspectives. Early works such as [zhang2013optical, jun2018using, 8460583, 9312948, ozguner2020camera] proposed external sensing, including optical tracking and depth/vision systems, to compensate for kinematic error. However, these approaches require complex physical setups and are difficult to replicate or deploy in new environments. More recently, vision-based frameworks such as [li2020super, Super2023, lu2021superdeepsurgicalperception] have achieved markerless tool tracking by utilizing keypoint detection and a Perspective-n-Point (PnP) solver. [Florian_filter] infused robot kinematic readings with 2D image-based feature detection to jointly estimate surgical instrument pose and kinematic calibration, formulating it as a lumped error. Nevertheless, vision-based feature detection, particularly keypoint detection, remains susceptible to noise and degradation under suboptimal endoscopic imaging conditions, making reliable pose initialization and stable tracking challenging. Liang et al. [liang2025differentiablerenderingbasedposeestimation] first incorporated differentiable rendering into surgical instrument pose reconstruction as a more robust alternative to PnP-based approaches. Rendering-based frameworks [liang2025differentiablerenderingbasedposeestimation, SurgiPose, yang2025instrumentsplattingcontrollablephotorealisticreconstruction] improve robustness and consistency by fully leveraging differentiable rendering matching in an iterative optimization manner, making them ideal for both single-frame pose reconstruction and kinematic calibration. However, they are computationally slow due to the iterative nature and prone to poor convergence, which limits their applicability in tracking setups with strict time constraints.
@@
To address these limitations and fully leverage the potential of rendering-based approaches for real-time tracking under noisy kinematic measurements, we propose a novel framework based on evolutionary optimization that jointly estimates surgical instrument pose and joint angles. By leveraging the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [Hansen2001-tj], which evaluates multiple pose candidates in parallel (Fig. 1), the proposed method substantially reduces per-frame iterations and mitigates convergence to local minima, thereby outperforming differentiable rendering. Compared to prior online tracking methods, this framework achieves significant improvements in both accuracy and runtime with GPU-accelerated batch rendering and fitness evaluation. Furthermore, the framework is extended to handle joint angle-free and bi-manual setups, enabling the calibration of online surgical videos for surgical robot learning applications using only visual inputs.
@@ II Related Work
Vision-based frameworks for robotic surgical tool tracking typically leverage deep neural networks to detect semantic visual features on the instrument and recover its pose using a PnP solver [8981600, 8932886]. However, for surgical robots such as da Vinci Research Kit (dVRK) [6907809], pose reconstruction must also account for inaccuracies in joint angle measurements induced by cable-driven transmission. To address this issue, Richter et al. [Florian_filter] propose a unified framework that incorporates keypoint detections as observations while estimating a lumped error transformation, enabling simultaneous estimation of the pose and kinematic errors. Building upon this formulation, [10610378] further improves the observation model with deep learning-based cylinder detection. However, these approaches are often unreliable due to the poor endoscopic video quality. They are also limited by the scarcity of data with ground truth annotations.
More recently, rendering-based approaches have attracted increasing attention due to their consistency and robustness achieved through direct silhouette matching. Liang et al. [liang2025differentiablerenderingbasedposeestimation] integrate geometric primitives with differentiable rendering-based optimization to enable robust single-shot calibration. Chen et al. [SurgiPose] and Yang et al. [yang2025instrumentsplattingcontrollablephotorealisticreconstruction] further apply differentiable rendering to recover surgical instrument trajectories from monocular videos. However, gradient-based optimization in these frameworks remains prone to local minima and slow convergence. Although coarse sampling [liang2025differentiablerenderingbasedposeestimation, SurgiPose] and image matching [yang2025instrumentsplattingcontrollablephotorealisticreconstruction] strategies have been introduced to improve initialization, the overall optimization process is still computationally expensive. Consequently, existing rendering-based approaches are primarily deployed in offline settings.
@@ III Methodology
An overview of the proposed pipeline is illustrated in Fig. 2, which estimates the 6-DoF pose and the visible joint angles of the surgical instrument with respect to the camera frame. Given RGB video frames, instrument segmentation masks are generated in real time using Surgical SAM 2 [liu2024surgicalsam2realtime], while a learned keypoint detection network localizes the tool tips. These visual observations define a unified optimization objective over pose and joint parameters. The resulting optimization problem is solved using CMA-ES [Hansen2001-tj], an evolution strategy that iteratively refines a Gaussian search distribution over candidate states, with efficient objective evaluations carried out through batch rendering. A Kalman filter with a constant-velocity motion model is incorporated to enforce temporal consistency, and the filtered state is used to initialize the search distribution for the subsequent frame.
@@ III-A Surgical Instrument Tracking
This study considers the Patient Side Manipulator (PSM) equipped with a Large Needle Driver (LND) from the da Vinci® Surgical System. The manipulator has 6 DoF, along with an additional gripper angle. As shown in Fig. 2, the end-effector coordinate frame is defined at the wrist pitch frame (joint 4) of the PSM. Similar to prior works [DR_orig, liang2025differentiablerenderingbasedposeestimation], the surgical instrument localization is formulated as a nonlinear optimization problem over the end-effector pose $\mathbf{T}_{r}^{c}\in SE(3)$, together with the last three visible joint angles $\mathbf{q}=[q_{1},q_{2},q_{3}]^{\top}\in\mathbb{R}^{3}$ corresponding to wrist pitch, wrist yaw, and gripper angle. For each RGB video frame $\mathbb{I}\in\mathbb{R}^{H\times W\times 3}$, the objective
@@ III-A1 Evolutionary Optimization
Rather than relying on gradient-based optimization, the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [Hansen2001-tj] is adopted. CMA-ES is a blackbox optimization method that iteratively refines a search distribution over the state space by maximizing the expected fitness $J(\mu,\Sigma):=\mathbb{E}\big[-\mathcal{L}(\mathbb{I},\Theta)\mid\mu,\Sigma\big]$, leading to the optimal parameters
@@
where $\mu,\Sigma$ defines a Gaussian distribution over the state variable $\Theta$. This formulation effectively lifts the original objective $\mathcal{L}(\mathbf{I},\Theta)$ from the parameter level to the distribution level, such that the loss landscape with respect to the mean $\mu$ approximates a Gaussian-smoothed version of $\mathcal{L}(\mathbf{I},\Theta)$, improving robustness to local minima; and covariance adaptations enable rapid convergence to locally quadratic optima of the expected fitness $J(\mu,\Sigma)$ [XNES].
@@ III-A2 Loss Function
where the rendering loss includes the Mean Squared Error (MSE) loss and the appearance loss [DR_orig, liang2025differentiablerenderingbasedposeestimation] to adjust the distance of the robot model
@@
The segmentation $\mathbb{M}_{\text{ref}}$ is produced by Surgical SAM 2 [liu2024surgicalsam2realtime] given the RGB image $\mathbb{I}$. The rendered mask $\mathbb{S}\in\{0,1\}^{H\times W}$ of the surgical tool is obtained from a blackbox renderer $\mathcal{R}(\cdot)$ given the mesh vertices transformed to the camera frame
@@ III-A3 State Parameterization
A unified rotation representation introduced by Liang et al. [liang2025differentiablerenderingbasedposeestimation] is adopted for pose sampling within CMA-ES. Specifically, the look-at camera representation $\alpha,\beta,\gamma]^{\top}\in\mathbb{R}^{3}$ is used, such that
@@
In addition, joint angles are subject to physical box constraints, while CMA-ES does not natively support constrained optimization. Hard clamping or sigmoid-based mappings, although commonly used in gradient-based methods, can distort the sampling distribution and introduce bias toward boundary values when applied within the evolution strategy [caraffini2019infeasibility]. A cosine-based reparameterization is adopted to enforce the constraints while maintaining a well-behaved search distribution. Denoting the lower and upper bounds of the $i$-th joint angle as $q^{\text{lb}}_{i}$ and $q^{\text{ub}}_{i}$, the mapping from a valid joint angle reading $q_{i}\in[q^{\text{lb}}_{i},q^{\text{ub}}_{i}]$ to an unconstrained representation $\hat{q}_{i}$ for CMA-ES optimization is given by
@@ III-B Bi-Manual Tracking
Inspired by separable CMA-ES [sepCMAES], we constrain the covariance update (4) in CMA-ES to a block-diagonal form:
@@ IV Experiments
All experiments are conducted on a workstation equipped with an NVIDIA RTX 4090 GPU and an AMD Ryzen 9 9960X CPU. Reference segmentation masks used in the tracking pipeline are generated by Surgical SAM 2 [liu2024surgicalsam2realtime] with point prompts provided in the first frame of the trajectory. The optimization frameworks are implemented with EvoTorch [evotorch] and NvDiffRast [Laine2020diffrast]. For evolutionary optimization, gradients are not evaluated, and instance-mode batch rendering is used for GPU-parallelized evaluation of the generated samples in each iteration.
For initialization, the pose-ranking method of Liang et al. [liang2025differentiablerenderingbasedposeestimation] is adapted to enable a fully joint angle-free pipeline. In addition to the original pose hypothesis space, initial joint configurations are also sampled from a feasible set. Instead of uniformly sampling this expanded search space, TuRBO [eriksson2019scalable] is applied to adaptively concentrate exploration on promising regions, thereby improving efficiency despite the increased dimensionality introduced by the additional joint variables.
During tracking, the parameter space is normalized to account for heterogeneous units and magnitudes. CMA-ES is initialized with covariance $\Sigma^{(0)}=\mathbf{I}$, and the default learning rates $c_{m}$ and $c_{\mu}$ are increased to accelerate convergence, with $N=70$ samples generated per iteration. The Adam optimizer [kingma2017adammethodstochasticoptimization] with a learning rate of $0.1$ in the normalized space serves as a gradient-based baseline.
@@
For real-world evaluation, 12 trajectories from the SurgPose dataset [wu2025surgposedatasetarticulatedrobotic] are selected, with both manipulators equipped with Large Needle Drivers (LND). The dataset is divided into 8 sequences with nominal speed and 4 sequences exhibiting rapid movements. In addition, we collected 6 trajectories of synchronized left and right stereo image streams at $640\times 480$ with painted markers and joint angle readings to benchmark against the online tool tracking method proposed by Richter et al. [Florian_filter], which relies on colored markers on the instrument for keypoint detection and applies a particle filter to estimate the lumped error transform. The collected dataset also includes refined segmentation masks to enable more accurate quantitative evaluation. The recorded trajectories vary in length from 250 to 554 frames, with an average of 436 frames per sequence. For the fast-motion trajectories in SurgPose and in the collected dataset, the process noises of the velocity term in the Kalman filter motion model (13) are correspondingly increased to better accommodate larger inter-frame velocity deviations.
@@
The qualitative tracking results are shown in Fig. 3 and Fig. 4, illustrating the accuracy and robustness of different tracking strategies on both the SurgPose dataset and the collected dataset. On the SurgPose dataset (Fig. 3), the proposed method is compared against the gradient-based baseline for bi-manual tracking. Gradient descent suffers from stagnation in local optima, which leads to error accumulation and visible drift over time. In contrast, CMA-ES demonstrates stronger robustness to early tracking errors (e.g., at Frame $250$) even without joint angle readings. Furthermore, the tracking becomes more accurate when joint angle readings are incorporated, producing consistent alignment between the estimated pose and image observations across frames. Figure 4 presents the comparison with the particle filter baseline of Richter et al. [Florian_filter] on the collected dataset. Using 1000 particles as in their original configuration, the particle filter shows noticeable misalignment at the tool tips. In contrast, the proposed method achieves more accurate alignment while maintaining smoother and more stable tracking.
@@
Richter et al. (1k particles) [Florian_filter] $0.2476\pm 0.0515$ $19.22\pm 0.45$
@@
Table II: Comparison of online tracking accuracy and inference speed between Richter et al. [Florian_filter] and the proposed method. All values are reported as $\text{mean}\pm\text{std}$. For the proposed method, the reported FPS includes both Surgical SAM 2 inference and subsequent CMA-ES optimization time.
Table I reports the quantitative results on the synthetic dataset and the SurgPose dataset. The proposed framework based on CMA-ES is compared against gradient descent (GD) baselines and XNES [XNES], a natural evolution strategy closely related to CMA-ES. In gradient-based ablations, NvDiffRast is replaced with PyTorch3D [ravi2020pytorch3d], as used in prior works [DR_orig, liang2025differentiablerenderingbasedposeestimation].
@@
Across both single- and dual-arm tracking, CMA-ES achieves the best balance between accuracy and efficiency. With only 3 iterations per frame, it consistently outperforms GD with 10–20 iterations in pose reconstruction for both the synthetic and real-world data while requiring approximately $37\%$ of the runtime. Incorporating joint angle readings further improves performance and yields the best overall results. Although XNES is more efficient than GD, it remains inferior to CMA-ES in reconstruction accuracy, as the CMA-ES update rule (4) enables faster convergence to local quadratic optima [XNES]. Substituting NvDiffRast with PyTorch3D in the gradient-based setting results in both reduced accuracy and increased runtime. Furthermore, the contribution of the keypoint loss $\mathcal{L}_{\text{kpts}}$ in (8) is assessed through two ablations: (1) removing the keypoint term slightly degrades performance, while (2) replacing the trained keypoint detection model with the traditional GoodFeaturesToTrack method [opencv] yields the worst overall performance, underscoring the importance of robust keypoint supervision.
Table II reports the quantitative results of benchmarking the proposed method against the particle filters proposed by Richter et al. [Florian_filter] for online tool tracking on the collected dataset. The mask error $(1-\text{IoU})$ and inference speed (FPS) are reported for both methods. For the proposed method, the FPS accounts for both Surgical SAM 2 prediction and CMA-ES optimization. The method demonstrates clear improvements in both accuracy and efficiency, benefiting from a GPU-parallelized implementation.

### 13 Mar 2026 v2

@@ I INTRODUCTION
Robot-assisted Minimally-Invasive Surgery (RMIS) has been increasingly adopted to improve surgical safety and procedural efficiency in complex, long-duration operations under physically demanding conditions. Accurate and efficient pose estimation of articulated surgical instruments is essential for a wide range of downstream applications, such as vision-based control [1], augmented reality guidance [2], and robot learning [3].
Uneven geometric structures, partial visibility of the kinematic chain, and noisy joint angle measurements caused by cable stretch introduce additional challenges for accurate articulated instrument pose recovery. Previous works have approached this problem from both kinematic calibration and pose calibration perspectives. Early works such as [4, 5, 6, 7, 8] proposed external sensing, including optical tracking and depth/vision systems, to compensate for kinematic error. However, these approaches require complex physical setups and are difficult to replicate or deploy in new environments. More recently, vision-based frameworks such as [9, 10, 11] have achieved markerless tool tracking by utilizing keypoint detection and a Perspective-n-Point (PnP) solver. [12] infused robot kinematic readings with 2D image-based feature detection to jointly estimate surgical instrument pose and kinematic calibration, formulating it as a lumped error. Nevertheless, vision-based feature detection, particularly keypoint detection, remains susceptible to noise and degradation under suboptimal endoscopic imaging conditions, making reliable pose initialization and stable tracking challenging. Liang et al. [13] first incorporated differentiable rendering into surgical instrument pose reconstruction as a more robust alternative to PnP-based approaches. Rendering-based frameworks [13, 14, 15] improve robustness and consistency by fully leveraging differentiable rendering matching in an iterative optimization manner, making them ideal for both single-frame pose reconstruction and kinematic calibration. However, they are computationally slow due to the iterative nature and prone to poor convergence, which limits their applicability in tracking setups with strict time constraints.
@@
To address these limitations and fully leverage the potential of rendering-based approaches for real-time tracking under noisy kinematic measurements, we propose a novel framework based on evolutionary optimization that jointly estimates surgical instrument pose and joint angles. By leveraging the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [16], which evaluates multiple pose candidates in parallel (Fig. 1), the proposed method substantially reduces per-frame iterations and mitigates convergence to local minima, thereby outperforming differentiable rendering. Compared to prior online tracking methods, this framework achieves significant improvements in both accuracy and runtime with GPU-accelerated batch rendering and fitness evaluation. Furthermore, the framework is extended to handle joint angle-free and bi-manual setups, enabling the calibration of online surgical videos for surgical robot learning applications using only visual inputs.
@@ II Related Work
Vision-based frameworks for robotic surgical tool tracking typically leverage deep neural networks to detect semantic visual features on the instrument and recover its pose using a PnP solver [21, 22]. However, for surgical robots such as da Vinci Research Kit (dVRK) [23], pose reconstruction must also account for inaccuracies in joint angle measurements induced by cable-driven transmission. To address this issue, Richter et al. [12] propose a unified framework that incorporates keypoint detections as observations while estimating a lumped error transformation, enabling simultaneous estimation of the pose and kinematic errors. Building upon this formulation, [24] further improves the observation model with deep learning-based cylinder detection. However, these approaches are often unreliable due to the poor endoscopic video quality. They are also limited by the scarcity of data with ground truth annotations.
More recently, rendering-based approaches have attracted increasing attention due to their consistency and robustness achieved through direct silhouette matching. Liang et al. [13] integrate geometric primitives with differentiable rendering-based optimization to enable robust single-shot calibration. Chen et al. [14] and Yang et al. [15] further apply differentiable rendering to recover surgical instrument trajectories from monocular videos. However, gradient-based optimization in these frameworks remains prone to local minima and slow convergence. Although coarse sampling [13, 14] and image matching [15] strategies have been introduced to improve initialization, the overall optimization process is still computationally expensive. Consequently, existing rendering-based approaches are primarily deployed in offline settings.
@@ III Methodology
An overview of the proposed pipeline is illustrated in Fig. 2, which estimates the 6-DoF pose and the visible joint angles of the surgical instrument with respect to the camera frame. Given RGB video frames, instrument segmentation masks are generated in real time using Surgical SAM 2 [25], while a learned keypoint detection network localizes the tool tips. These visual observations define a unified optimization objective over pose and joint parameters. The resulting optimization problem is solved using CMA-ES [16], an evolution strategy that iteratively refines a Gaussian search distribution over candidate states, with efficient objective evaluations carried out through batch rendering. A Kalman filter with a constant-velocity motion model is incorporated to enforce temporal consistency, and the filtered state is used to initialize the search distribution for the subsequent frame.
@@ III-A Surgical Instrument Tracking
This study considers the Patient Side Manipulator (PSM) equipped with a Large Needle Driver (LND) from the da Vinci® Surgical System. The manipulator has 6 DoF, along with an additional gripper angle. As shown in Fig. 2, the end-effector coordinate frame is defined at the wrist pitch frame (joint 4) of the PSM. Similar to prior works [26, 13], the surgical instrument localization is formulated as a nonlinear optimization problem over the end-effector pose $\mathbf{T}_{r}^{c}\in SE(3)$, together with the last three visible joint angles $\mathbf{q}=[q_{1},q_{2},q_{3}]^{\top}\in\mathbb{R}^{3}$ corresponding to wrist pitch, wrist yaw, and gripper angle. For each RGB video frame $\mathbb{I}\in\mathbb{R}^{H\times W\times 3}$, the objective
@@ III-A1 Evolutionary Optimization
Rather than relying on gradient-based optimization, the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [16] is adopted. CMA-ES is a blackbox optimization method that iteratively refines a search distribution over the state space by maximizing the expected fitness $J(\mu,\Sigma):=\mathbb{E}\big[-\mathcal{L}(\mathbb{I},\Theta)\mid\mu,\Sigma\big]$, leading to the optimal parameters
@@
where $\mu,\Sigma$ defines a Gaussian distribution over the state variable $\Theta$. This formulation effectively lifts the original objective $\mathcal{L}(\mathbf{I},\Theta)$ from the parameter level to the distribution level, such that the loss landscape with respect to the mean $\mu$ approximates a Gaussian-smoothed version of $\mathcal{L}(\mathbf{I},\Theta)$, improving robustness to local minima; and covariance adaptations enable rapid convergence to locally quadratic optima of the expected fitness $J(\mu,\Sigma)$ [27].
@@ III-A2 Loss Function
where the rendering loss includes the Mean Squared Error (MSE) loss and the appearance loss [26, 13] to adjust the distance of the robot model
@@
The segmentation $\mathbb{M}_{\text{ref}}$ is produced by Surgical SAM 2 [25] given the RGB image $\mathbb{I}$. The rendered mask $\mathbb{S}\in\{0,1\}^{H\times W}$ of the surgical tool is obtained from a blackbox renderer $\mathcal{R}(\cdot)$ given the mesh vertices transformed to the camera frame
@@ III-A3 State Parameterization
A unified rotation representation introduced by Liang et al. [13] is adopted for pose sampling within CMA-ES. Specifically, the look-at camera representation $\alpha,\beta,\gamma]^{\top}\in\mathbb{R}^{3}$ is used, such that
@@
In addition, joint angles are subject to physical box constraints, while CMA-ES does not natively support constrained optimization. Hard clamping or sigmoid-based mappings, although commonly used in gradient-based methods, can distort the sampling distribution and introduce bias toward boundary values when applied within the evolution strategy [28]. A cosine-based reparameterization is adopted to enforce the constraints while maintaining a well-behaved search distribution. Denoting the lower and upper bounds of the $i$-th joint angle as $q^{\text{lb}}_{i}$ and $q^{\text{ub}}_{i}$, the mapping from a valid joint angle reading $q_{i}\in[q^{\text{lb}}_{i},q^{\text{ub}}_{i}]$ to an unconstrained representation $\hat{q}_{i}$ for CMA-ES optimization is given by
@@ III-B Bi-Manual Tracking
Inspired by separable CMA-ES [29], we constrain the covariance update (4) in CMA-ES to a block-diagonal form:
@@ IV Experiments
All experiments are conducted on a workstation equipped with an NVIDIA RTX 4090 GPU and an AMD Ryzen 9 9960X CPU. Reference segmentation masks used in the tracking pipeline are generated by Surgical SAM 2 [25] with point prompts provided in the first frame of the trajectory. The optimization frameworks are implemented with EvoTorch [30] and NvDiffRast [31]. For evolutionary optimization, gradients are not evaluated, and instance-mode batch rendering is used for GPU-parallelized evaluation of the generated samples in each iteration.
For initialization, the pose-ranking method of Liang et al. [13] is adapted to enable a fully joint angle-free pipeline. In addition to the original pose hypothesis space, initial joint configurations are also sampled from a feasible set. Instead of uniformly sampling this expanded search space, TuRBO [32] is applied to adaptively concentrate exploration on promising regions, thereby improving efficiency despite the increased dimensionality introduced by the additional joint variables.
During tracking, the parameter space is normalized to account for heterogeneous units and magnitudes. CMA-ES is initialized with covariance $\Sigma^{(0)}=\mathbf{I}$, and the default learning rates $c_{m}$ and $c_{\mu}$ are increased to accelerate convergence, with $N=70$ samples generated per iteration. The Adam optimizer [33] with a learning rate of $0.1$ in the normalized space serves as a gradient-based baseline.
@@
For real-world evaluation, 12 trajectories from the SurgPose dataset [34] are selected, with both manipulators equipped with Large Needle Drivers (LND). The dataset is divided into 8 sequences with nominal speed and 4 sequences exhibiting rapid movements. In addition, we collected 6 trajectories of synchronized left and right stereo image streams at $640\times 480$ with painted markers and joint angle readings to benchmark against the online tool tracking method proposed by Richter et al. [12], which relies on colored markers on the instrument for keypoint detection and applies a particle filter to estimate the lumped error transform. The collected dataset also includes refined segmentation masks to enable more accurate quantitative evaluation. The recorded trajectories vary in length from 250 to 554 frames, with an average of 436 frames per sequence. For the fast-motion trajectories in SurgPose and in the collected dataset, the process noises of the velocity term in the Kalman filter motion model (13) are correspondingly increased to better accommodate larger inter-frame velocity deviations.
@@
The qualitative tracking results are shown in Fig. 3 and Fig. 4, illustrating the accuracy and robustness of different tracking strategies on both the SurgPose dataset and the collected dataset. On the SurgPose dataset (Fig. 3), the proposed method is compared against the gradient-based baseline for bi-manual tracking. Gradient descent suffers from stagnation in local optima, which leads to error accumulation and visible drift over time. In contrast, CMA-ES demonstrates stronger robustness to early tracking errors (e.g., at Frame $250$) even without joint angle readings. Furthermore, the tracking becomes more accurate when joint angle readings are incorporated, producing consistent alignment between the estimated pose and image observations across frames. Figure 4 presents the comparison with the particle filter baseline of Richter et al. [12] on the collected dataset. Using 1000 particles as in their original configuration, the particle filter shows noticeable misalignment at the tool tips. In contrast, the proposed method achieves more accurate alignment while maintaining smoother and more stable tracking.
@@
Richter et al. (1k particles) [12] $0.2476\pm 0.0515$ $19.22\pm 0.45$
@@
Table II: Comparison of online tracking accuracy and inference speed between Richter et al. [12] and the proposed method. All values are reported as $\text{mean}\pm\text{std}$. For the proposed method, the reported FPS includes both Surgical SAM 2 inference and subsequent CMA-ES optimization time.
Table I reports the quantitative results on the synthetic dataset and the SurgPose dataset. The proposed framework based on CMA-ES is compared against gradient descent (GD) baselines and XNES [27], a natural evolution strategy closely related to CMA-ES. In gradient-based ablations, NvDiffRast is replaced with PyTorch3D [35], as used in prior works [26, 13].
@@
Across both single- and dual-arm tracking, CMA-ES achieves the best balance between accuracy and efficiency. With only 3 iterations per frame, it consistently outperforms GD with 10–20 iterations in pose reconstruction for both the synthetic and real-world data while requiring approximately $37\%$ of the runtime. Incorporating joint angle readings further improves performance and yields the best overall results. Although XNES is more efficient than GD, it remains inferior to CMA-ES in reconstruction accuracy, as the CMA-ES update rule (4) enables faster convergence to local quadratic optima [27]. Substituting NvDiffRast with PyTorch3D in the gradient-based setting results in both reduced accuracy and increased runtime. Furthermore, the contribution of the keypoint loss $\mathcal{L}_{\text{kpts}}$ in (8) is assessed through two ablations: (1) removing the keypoint term slightly degrades performance, while (2) replacing the trained keypoint detection model with the traditional GoodFeaturesToTrack method [36] yields the worst overall performance, underscoring the importance of robust keypoint supervision.
Table II reports the quantitative results of benchmarking the proposed method against the particle filters proposed by Richter et al. [12] for online tool tracking on the collected dataset. The mask error $(1-\text{IoU})$ and inference speed (FPS) are reported for both methods. For the proposed method, the FPS accounts for both Surgical SAM 2 prediction and CMA-ES optimization. The method demonstrates clear improvements in both accuracy and efficiency, benefiting from a GPU-parallelized implementation.